OpenStudy (anonymous):

(sinx)(cotx+cosxtanx)=cosx+sin^2x

OpenStudy (anonymous):

\[\sin x(\cot x+\cos x\tan x)\]\[=\sin x\left(\frac{\cos x}{\sin x}+\cos x \frac{\sin x}{\cos x}\right)\] [Then cancellation and distribution.]

OpenStudy (anonymous):

how do i cancel and distribute?

OpenStudy (anonymous):

You can multiply \(\sin x\) to \(each\) of \(\frac{\cos x}{\sin x}\) and \(\cos x \frac{\sin x}{\cos x}\).

OpenStudy (anonymous):

what happens with cos x?

OpenStudy (anonymous):

cos x of the first expression remains, since only the sin x cancels, while the cos x of the second expression immediately cancels.

OpenStudy (anonymous):

Like this:\[\sin x \left(\frac{\cos x}{\sin x}+\cos x \frac{\sin x}{\cos x} \right)\]\[=\sin x \frac{\cos x}{\sin x}+\sin x \cos x \frac{\sin x}{\cos x} \]\[=\cos x+\sin x\sin x\]

OpenStudy (anonymous):

okay got it. thanks!